OpenStudy (anonymous):

the intigral from 0 to pi/2 of (e^(2-cosx))sinx dx =?

OpenStudy (bahrom7893):

Let u = 2-Cos(x)

OpenStudy (bahrom7893):

du = Sinx dx

OpenStudy (bahrom7893):

Integral is from 0 to pi/2 but now I said let u = 2-Cos(x), so: Lower limit: u(o) = 2 - Cos(0) = 2 - 1 = 1

OpenStudy (bahrom7893):

Upper limit: u(pi/2) = 2 - Cos(pi/2) = 2-0 = 2

OpenStudy (bahrom7893):

So now you have: Integral from 1 to 2 e^u * du = e^u (from 1 to 2)

OpenStudy (bahrom7893):

e^2 - e = e(e-1)

OpenStudy (anonymous):

ok tyvm that was a big help

OpenStudy (bahrom7893):

Please click on Become a fan, thanks! =)

OpenStudy (anonymous):

already am lol

OpenStudy (bahrom7893):

lol thanks! =)

OpenStudy (anonymous):

ty your the one that has answered my questions